OpenStudy (anonymous):

I can't seem to see it... Is there a way to simplify (9-x^2)/(x-3) ?

jhonyy9 (jhonyy9):

us a2 -b2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

(9-x2)=-(x2-9)=-(x-3)(x+3) after can be simplify x-3 and will remain -(x+3) what is equal 3-x

jhonyy9 (jhonyy9):

the last part is wrong sorry will remain at finaly -x-3

jhonyy9 (jhonyy9):

or -(x+3)

jhonyy9 (jhonyy9):

how do you like to write it

jhonyy9 (jhonyy9):

Denebel is it understandably sure ?

jhonyy9 (jhonyy9):

hey ! are you there ?

jhonyy9 (jhonyy9):

i like to learn you

jhonyy9 (jhonyy9):

because this is very very easy

OpenStudy (anonymous):

Oh thank you :) why couldn't I see that earlier.. haha